# A Homothetic Gauge Theory

### v2

@@ Abstract
While the principle of local gauge invariance forms the bedrock of modern physics, the symmetry of scale invariance has traditionally been treated on a different footing. This paper develops a homothetic gauge theory (HGT) that promotes local scale transformations to a genuine gauge symmetry, placing them within the same geometric framework as conventional interactions. We introduce a novel mathematical structure based on an extended vector bundle that linearizes the action of local scaling, allowing for the definition of a homothetic connection and a corresponding gauge-covariant field strength. A central mathematical innovation of this theory is the development of a homothetic Hodge–de Rham complex and cohomology, wherein the standard differential and codifferential operators are deformed in a scale-covariant manner. When this formalism is applied to classical electromagnetism, as a U(1) gauge theory, the generalized homothetic Maxwell equations are obtained, and a direct connection is established with the spontaneous symmetry breaking. As a significant physical application, this framework provides a compelling resolution to the classical problem of the infinite self-energy of a point charge. By modeling the source as a penalized Dirichlet boundary condition, we construct a non-singular electrostatic potential.
Keywords: gauge theory, scale invariance, local symmetries, vector bundles, field theory, Weyl integrable geometry, Hodge–de Rham cohomology, spontaneous symmetry breaking, Maxwell equations, potential theory
@@ 1 Introduction
At the heart of modern theoretical physics lies the principle of local gauge invariance, a powerful idea that dictates the fundamental interactions of nature by elevating global symmetries to local, spacetime-dependent ones [8]. One of the most well-established gauge theories in physics is classical electromagnetism, formulated as a $U(1)$ gauge theory. Its mathematical foundation rests on the geometry of principal bundles and connections, offering a prototypical example of how gauge symmetry can encode physical interactions [14]. By extending this geometric framework to incorporate local scale transformations—homotheties—we aim to generalize the notion of gauge invariance itself. Such a generalization offers novel geometric and conceptual reinterpretations of gauge structure, aligning with recent insights into the foundations of gauge symmetries and their reformulations [10].
The paper establishes a new homothetic gauge theory where the gauge freedom corresponds to local changes of scale. The central innovation is the linear structure endowed upon scale transformations, which naturally integrates into the standard fiber bundle formalism. It occurs by embedding the tangent and cotangent bundles in the doubled bundles, anchored by a fixed but arbitrary reference vector field. It transforms the non-linear action of scaling into a linear matrix operation on the fibers of these extended bundles.
@@
In the paper, the Weyl integrable geometry is obtained as a special case of the general formulation, and it is shown that a homothetic cohomology is definable consistent with this very conformal geometry, resulting in a generalized de Rham complex. When applied to the context of electromagnetism, the theory yields a generalization of Maxwell’s equations. These ”homothetic Maxwell equations” feature new interaction terms coupling the electromagnetic field to the gradient of the scalar gauge field that parameterizes the local scale.
@@
A feature of HGT is that it clarifies the usual split between symmetry of the equations and symmetry of their solutions [10, 25]. Instead of purely multiplicative Weyl rescalings, we employ an affine action generated by the dilatation one-form $b=d\lambda$ with offset fields $\phi_{d}$. The field equations remain homothetically covariant, while physically meaningful scales enter through initial/boundary data and gauge choices. In this manner, “spontaneous” breaking of local scale invariance arises from selecting a background/gauge (fixing $\lambda$) and offsets $\phi_{d}$ that act as order parameters, not from a Higgs-type potential. The equations keep homothetic covariance; particular solution classes pick a preferred scale via the imposed data (see $\S$ 7) [33, 25].
The generalized Maxwell sector (Secs. 5.1 &\& 7.1) makes the above issue concrete: the homothetic Lorenz gauge and associated wave/Laplace operators support a penalization that enforces Dirichlet data on a shrinking interface. In the singular limit one recovers the exterior electrovac and avoids the point-charge self-energy divergence; the emergent scale is set by the penalization profile and choice of $\lambda$ (Sec. 8.2), so symmetry is preserved in the equations but broken by solution selection.
@@
The paper is structured as follows. Section 2 introduces the block–homothetic gauge group and the doubled-bundle construction that renders local scale transformations linear. Section 3 embeds this algebraic setup in a generalized Weyl-integrable geometry, deriving the compatible connection and curvature. In Section 4 we develop the homothetic de Rham complex and prove an extended Hodge decomposition that underpins the subsequent field theory. Section 5 specializes the formalism to electromagnetism, obtaining homothetically covariant Maxwell equations and a modified Lorentz force. Section 6 then derives the associated wave and Laplace equations. Section 7 discusses the relation between HGT and spontaneous symmetry breaking in U(1)-symmetric theories. The physical and mathematical applications of the theory are provided in two related sections. Section 8.1 analyses the Cauchy, Dirichlet, and Neumann boundary-value problems; the discussions that are applied in Section 8.2 to construct a non-singular electrostatic potential for a point charge. Section 9 summarises the main findings. For the sake of continuity, some technical materials are gathered in five appendices at the end of the paper.
@@ 2 First setup: Block-homothetic Transformations
The above construction, while self-contained, is not arbitrary. As we demonstrate in Appendix A, this entire framework arises naturally as a specific gauge fixing within the standard formulation of conformal Cartan geometry, providing a deeper geometric justification for our approach. Moreover, in appendix C the connection and curvature of $\mathcal{H}$ bundle, as a principal bundle, are obtained.
@@ 3 A Generalized Weyl Integrable Geometry
It should be stressed that although we finally set $X_{d}=0$, this is not just a complicated way to achieve the standard conformal geometry. In fact, we will always consider this geometry as a homothetic geometry with the zero homothety center. In particular, it manifests itself in the differential forms we will associate with it in the next section— see Eq. (20) and the discussions after that.
On the other hand, nothing in our formulation prevents us from admitting the conventional Weyl structure. Recall that [30] a Weyl integrable geometry is defined by the triple $(M,[\eta],\phi)$, where $\eta]$ is the conformal class of Lorentzian metrics and $\phi:=d\lambda$ is a globally defined, closed 1-form satisfying
@@ 4.1 The Homothetically Transformed Differential Forms
obtainable from the first row of $\hat{\alpha}=\Lambda\bar{\alpha}$. Associating such a transformation to the conformally transformed spacetime is mathematically legitimated (see appendix A and the discussion below). Moreover, the following special cases are noticeable:
@@ 4.2 Generalized de Rham complex and Hodge decomposition
in which the space of homothetically transformed $k$-forms splits orthogonally into exact, coexact, and harmonic components, all defined relative to the Weyl-rescaled metric $\tilde{\eta}=e^{-2\lambda(x)}\eta$. This decomposition underlies the forthcoming formulation of field equations—most notably, generalized Maxwell-type theories—in which the harmonic representatives encode gauge-invariant field configurations within the homothetic background. A functional analytic proof of decomposition (31) is provided in appendix D.
@@ 5 Homothetic Electromagnetism
The above homothetic Hodge-de Rham complex can be constructed for the electromagnetism, as a $U(1)$ gauge theory defined on the Minkowski spacetime manifold. It means embedding any $k$-form $\alpha^{k}$, where $0\leq k\leq 4$ in $\Omega^{k}(M)\oplus\Omega^{k}(M)$, where $\mathcal{H}$ can act on, making $\hat{\Omega}^{k}$. However, in order for the homothetic transformations to preserve the symmetries of the electromagnetism, they must act directly on the connections of the principal bundle, that is, the gauge potential fields $A=A_{\mu}dx^{\mu}$ (as the $\mathfrak{u}(1)$-valued 1-forms on $M$). Such a setup results in Fig. 3, in which we defined:
$\bar{A}:=\begin{pmatrix}A\\ A_{d}\end{pmatrix},\qquad\hat{A}:=\begin{pmatrix}\tilde{A}\\ A_{d}\end{pmatrix}=\Lambda\bar{A},$ (32)
and the other $k\neq 1$ forms are transformed with the offset forms $\delta A_{d}$, $dA_{d}$, $d^{2}A_{d}=0$, etc., inducing by this genuine transformation. For example, by definition of:
$\bar{F}:=\begin{pmatrix}F\\ F_{d}\end{pmatrix},\qquad\hat{F}:=\begin{pmatrix}\tilde{F}\\ F_{d}\end{pmatrix}=\Lambda\bar{F},$ (33)
the offset is $F_{d}:=dA_{d}$, where $A_{d}$ has already defined.
5.1 Sourceless Maxwell Equations
On the 4-dimensional Minkowski spacetime $(M,\eta)$, classical (source-free) electromagnetism is expressed as:
$dF=0,$ (34)
$\delta F=0,$ (35)
with $F=dA$ for some 1-form $A$ defined on $M$. We now consider an offset potential $A_{d}\in\Omega^{1}(M)$ with $F_{d}=dA_{d}$, and define the block forms $\hat{A}$ and $\hat{F}$ as above, and the Maxwell equations in the homothetically transformed geometry obtains as:
$\hat{d}\hat{F}=0,$ (36)
$\hat{\delta}\hat{F}=0.$ (37)
Then, $\hat{F}=\hat{d}\hat{A}$ is invariant under the homothetic gauge transformation:
$\boxed{\hat{A}\mapsto\hat{A}+\hat{d}\hat{\phi}}\qquad\text{where}\qquad\hat{\phi}=\Lambda\bar{\phi}=\Lambda\begin{pmatrix}\phi\\ \phi_{d}\end{pmatrix}$, (38)
and $\phi,\phi_{d}$ are arbitrary smooth functions on $M$. This is the homothetic equivalent of the classical gauge invariance of the Maxwell equations.
The first row of the homothetic Maxwell equations (36)-(37) yields:
$d\tilde{F}+(d\lambda)\wedge(\tilde{F}-F_{d})=0,$ (39)
$\delta\tilde{F}-*(d\lambda)\wedge*(\tilde{F}-F_{d})=0.$ (40)
As one can see, some additional terms have appeared which represent the interaction of the electromagnetic fields with the offset fields. Apparently, for $d\lambda\equiv 0$, these additional terms disappear, and the classical Maxwell equations retrieve.
@@ 5.1 Sourceless Maxwell Equations
By translation of the above equations on the tangent bundle, the generalized Maxwell equations are obtained in the more conventional three-vector formulation:
$\nabla\cdot\tilde{E}+\nabla\lambda\cdot(\tilde{E}-E_{d})$ $=0,\qquad\text{generalized Gauss Law};$ (41)
$\partial_{t}\tilde{E}-\nabla\times\tilde{B}+\partial_{t}\lambda(\tilde{E}-E_{d})-\nabla\lambda\times(\tilde{B}-B_{d})$ $=0,\qquad\text{generalized Ampère's Law};$ (42)
$\partial_{t}\tilde{B}+\nabla\times\tilde{E}+\partial_{t}\lambda(\tilde{B}-B_{d})+\nabla\lambda\times(\tilde{E}-E_{d})$ $=0,\qquad\text{generalized Faraday Law};$ (43)
$\nabla\cdot\tilde{B}=0,\quad\text{for }\nabla\cdot B_{d}$ $=0,\qquad\text{Gauss's Law for Magnetism}.$ (44)
This formulation offers a homothetically covariant generalization of electromagnetism. A distinguishing feature of this system is that it is sourceless. In fact, the emergence of penalty terms, with geometric origins, enable us to derive the dynamics of electromagnetic fields without the need for source terms. This issue is crucial in construction of a non-singular Model for a Point Electric Charge in $\S$ 8.2. However, for the sake of completeness, a non-homogeneous formulation is also provided in $\S$ 7.

### v3

@@ Abstract
We introduce a Homothetic Hodge–de Rham (HHDR) theory that extends the de Rham complex to homothetically transformed differential forms and develops the associated homothetic Hodge machinery. Enforcing homothetic symmetry on physical laws naturally yields scale-covariant couplings, interpretable as canonical penalty terms in the framework of differential equations. Building on this foundation, we construct a Homothetic Gauge Theory (HGT) and apply it to classical electromagnetism, leading to generalized homothetic Maxwell equations. The approach establishes a link between homothetic scaling and conformal field theory concepts such as the dilaton field and vacuum expectation values (VEVs). As a concrete physical application, we address the classical divergence of point-charge self-energy: by modeling the source as a penalized Dirichlet boundary condition, we obtain a non-singular electrostatic potential. This framework highlights a mathematically rigorous and physically meaningful extension of gauge theory with potential relevance across field theory and electrodynamics.
journal: Annals of Physics
Keywords: Homothetic Gauge Theory, Hodge–de Rham Theory, Scale Symmetry, Conformal Field Theory (CFT), Classical Electrodynamics, Point Charge Self-Energ
@@ 1 Introduction
The principle of local gauge invariance, which dictates fundamental interactions by promoting global symmetries to local ones, is a cornerstone of modern physics [10]. Classical electromagnetism, a $U(1)$ gauge theory built on the geometry of principal bundles, provides the primary model for this mechanism [15]. This work generalizes the notion of gauge invariance itself by extending this framework to include local scale transformations, or homotheties. Our approach offers novel reinterpretations of gauge structure that align with modern investigations into the foundations of gauge symmetry [12].
The paper establishes a new homothetic gauge theory (HGT) where the gauge freedom corresponds to local changes of scale. The central innovation is the linear structure endowed upon scale transformations, which naturally integrates into the standard fiber bundle formalism. It occurs by embedding the bundles in the doubled bundles, anchored by a fixed but arbitrary reference vector field. It transforms the non-linear operation of scaling into a linear matrix operation on the fibers of these extended bundles.
@@
In the paper, the Weyl integrable geometry is obtained as a special case of a more general formulation. It is shown that a homothetic cohomology is definable, consistent with this very conformal geometry. It results in a generalized (homothetic) de Rham complex. When applied to the context of electromagnetism, the theory yields a generalization of Maxwell’s equations. These "homothetic Maxwell equations" feature new interaction terms coupling the electromagnetic field to the gradient of the scalar gauge field that parameterizes the local scale.
To connect this mathematical framework to established physics, we identify the scaling field $\lambda$ with the conformal field theory (CFT) dilaton, which acts as a compensator for local Weyl rescalings and is linked to the renormalization group flow. Its one-form, $d\lambda$, naturally plays the role of the Weyl connection in the field equations. Furthermore, we interpret the offset fields $\alpha_{d}$ as the vacuum expectation values (VEVs) of the gauge fields. In the weak-coupling limit, which corresponds to $\lambda\to\infty$, the theory dynamically selects a flat connection, ensuring that the offset fields represent the vacuum holonomy character of the theory.
@@
The paper is structured as follows. Section 2 introduces the block–homothetic gauge group and the doubled-bundle construction that renders local scale transformations linear. Section 3 embeds this algebraic setup in a generalized Weyl-integrable geometry, deriving the compatible connection and curvature. In Section 4 we develop the homothetic de Rham complex and prove an extended Hodge decomposition that underpins the subsequent field theory. Section 5 specializes the formalism to electromagnetism, obtaining homothetically covariant Maxwell equations and a modified Lorentz force. Section 6 then derives the associated wave and Laplace equations. In Section 7 we disentangle mathematical structure from physics by detailing the roles of the scaling field $\lambda$ and the offset fields $\alpha_{d}$ within HGT. The physical and mathematical applications of the theory are provided in two related sections. Section 8.1 analyses the Cauchy, Dirichlet, and Neumann boundary-value problems; the discussions that are applied in Section 8.2 to construct a non-singular electrostatic potential for a point charge. Section 9 summarises the main findings. For the sake of continuity, some technical materials are gathered in five appendices at the end of the paper.
@@ 2 First setup: Block-homothetic Transformations
The above construction, while self-contained, is not arbitrary. As we demonstrate in A, this entire framework arises naturally as a specific gauge fixing within the standard formulation of conformal Cartan geometry, providing a deeper geometric justification for our approach. Moreover, in C the connection and curvature of $\mathcal{H}$ bundle, as a principal bundle, are obtained.
@@ 3 A Generalized Weyl Integrable Geometry
It should be emphasized that although we ultimately set $X_{d}=0$, this is not merely a complicated way to achieve the standard conformal geometry. It leads us to the homothetic cohomology we will associate with it in the next sections. On the other hand, our spacetime manifold is endowed with the conformal structure, and nothing prevents us from considering a Weyl (integrable) connection (see A for the justification from the conformal Cartan geometry viewpoint).
Recall that [30] a Weyl integrable geometry is defined by the triple $(M,[\eta],\phi)$, where $\eta]$ is the conformal class of Lorentzian metrics and $\phi:=d\lambda$ is a globally defined, closed 1-form satisfying
@@ 4.1 The Homothetically Transformed Differential Forms
obtainable from the first row of $\hat{\alpha}=\Lambda\bar{\alpha}$. Associating such a transformation to the conformally transformed spacetime is mathematically legitimated (see A and the discussion below). Moreover, the following special cases are noticeable:
  • 1.
@@ 4.2 Generalized de Rham complex and Hodge decomposition
in which the space of homothetically transformed $k$-forms splits orthogonally into exact, coexact, and harmonic components, all defined relative to the Weyl-rescaled metric $\tilde{\eta}=e^{-2\lambda(x)}\eta$. This decomposition underlies the forthcoming formulation of field equations—most notably, generalized Maxwell-type theories—in which the harmonic representatives encode gauge-invariant field configurations within the homothetic background. A functional analytic proof of decomposition (31) is provided in D.
@@ 5 Homothetic Electromagnetism
Vacuum Maxwell theory is elegantly formulated using the de Rham complex and Hodge theory, a framework that naturally generalizes to HHDR theory. This generalization is achieved by embedding each $k$-form $\alpha^{k}$ (for $0\leq k\leq 4$) into the direct sum $\hat{\Omega}^{k}(M):=\Omega^{k}(M)\oplus\Omega^{k}(M)$, on which a homothetic operator $\mathcal{H}$ can act, for example:
$\bar{A}:=\begin{pmatrix}A\\ A_{d}\end{pmatrix},\qquad\hat{A}:=\begin{pmatrix}\tilde{A}\\ A_{d}\end{pmatrix}=\Lambda\bar{A},\qquad\bar{F}:=\begin{pmatrix}F\\ F_{d}\end{pmatrix},\qquad\hat{F}:=\begin{pmatrix}\tilde{F}\\ F_{d}\end{pmatrix}=\Lambda\bar{F},$ (32)
for the 1-form gauge field $A$, and the 2-form electromagnetic field $F$.
However, physically consistent determination of the offset fields still remains. Until working on the general HHDR, the offset fields are just mathematical objects. But when applying to electromagnetism, they must have physical meanings.
5.0.1 The Offset fields
In order to respect the $U(1)$ symmetry, the genuine homothetic transformation must be applied on the gauge fields $A$, that is, the $\mathfrak{u}(1)$-valued connection 1-forms (see Fig. 3), the fact that also support by the action formulation in the next section. Then the offset fields of the other ranks induce from $A_{d}$ as $\delta A_{d}$ for the 0-forms, and $dA_{d}$, $d^{2}A_{d}=0$, …for the 2-forms, 3-forms,…. Therefore, determination of the physically meaningful $A_{d}$ is in order.
We distinguish two different situations:
  1. 1.
    For the non-trivial topologies, or/and the presence of the boundaries or/and sources, physics do not put phenomenological restrictions on $A_{d}$. In this situation, $A_{d}$ and the other derived offset fields $\chi_{d}:=\delta A_{d}$ and $F_{d}:=dA_{d}$ are things that come from the boundaries, sources, and so on.
  2. 2.
    For the trivial topology, without any source and boundary, the offset fields have a clear connection with the vacuum expectation values (VEVs). The issue is discussed in detail in $\S$ 7.2. By identifying them as the VEVs, in order to have a Lorentz invariant vacuum, we must have:
    $\langle F_{\mu\nu}\rangle=0,$ (33)
    which means the trivial flat $A_{d}=\langle A\rangle$, which means in turn an exact $A_{d}$, that is, $A_{d}=d\chi$.
@@ 5.0.1 The Offset fields
5.1 Homothetic Maxwell equations
On Minkowski spacetime $(M,\eta)$ let $\Omega:=\Omega^{1}(M)/d\Omega^{0}(M)$ and $\bar{\Omega}:=\Omega\oplus\Omega$ with projection $\pi_{1}:\bar{\Omega}\to\Omega$. Using the previously defined homothetic map $\Lambda[\lambda]$ on 1-form doublets $\bar{A}$ and $\hat{A}$, we have
$A=A_{d}+e^{\lambda}(\tilde{A}-A_{d}).$
Starting from the genuine Maxwell functional on $(M,\eta)$,
$S_{\rm Max}[A;J]=-\frac{1}{2}\int F\wedge*_{\eta}F+\int A\wedge*_{\eta}J,\qquad F:=dA,\quad d(*_{\eta}J)=0,$ (34)
the HGT action is the pullback of $S_{\rm Max}$ along $\pi_{1}\circ\Lambda[\lambda]^{-1}$:
$\boxed{\ S_{\rm HGT}[\tilde{A},A_{d};\lambda]=\bigl(\pi_{1}\circ\Lambda[\lambda]^{-1}\bigr)^{*}S_{\rm Max}=S_{\rm Max}\!\left(A_{d}+e^{\lambda}(\tilde{A}-A_{d});\,J\right).}$ (35)
In $d=4$ the kinetic term is pointwise conformally invariant, so one may equivalently use $\tilde{\eta}=e^{-2\lambda}\eta$. With $\hat{d}$ and $\hat{\delta}$ as in $\S$ 4.2, the field-strength doublet is
$\hat{F}:=\hat{d}\,\hat{A}=\begin{pmatrix}\tilde{F}\\ F_{d}\end{pmatrix},\qquad\tilde{F}=d\tilde{A}+b\wedge(\tilde{A}-A_{d}),\quad F_{d}=dA_{d},\quad b:=d\lambda.$ (36)
Using the block pairing built from $_{\tilde{\eta}}$ on each slot, the HGT Maxwell action reads
$S_{\rm HGT}[\hat{A};\lambda;\hat{J}]=-\frac{1}{2}\int\!\big(\tilde{F}\wedge*_{\tilde{\eta}}\tilde{F}+F_{d}\wedge*_{\tilde{\eta}}F_{d}\big)+\int\!\big(\tilde{A}\wedge*_{\tilde{\eta}}\tilde{J}+A_{d}\wedge*_{\tilde{\eta}}J_{d}\big),$ (37)
with source doublet $\hat{J}=(\tilde{J},J_{d})^{\mathsf{T}}$ obeying $\hat{\delta}\,\hat{J}=0$ (upper slot couples to the physical current).
Varying (37) and integrating by parts with $\hat{\delta}$ gives the homothetic Maxwell equations:
$\boxed{\qquad\hat{\delta}\,\hat{F}=\hat{J},\qquad\hat{d}\,\hat{F}=0.\qquad}$ (38)
In components:
$\delta_{\tilde{\eta}}\tilde{F}+\iota_{b^{\sharp}}\tilde{F}$ $=\tilde{J}$, $d\tilde{F}+b\wedge\tilde{F}-b\wedge F_{d}$ $=0,$ (39)
$\delta_{\tilde{\eta}}F_{d}-\iota_{b^{\sharp}}\tilde{F}$ $=J_{d},$ $dF_{d}$ $=0.$ (40)
Setting $J=0$ (equivalently $\hat{J}=0$) in (38) yields the homogeneous system
$\hat{\delta}\,\hat{F}=0,\qquad\hat{d}\,\hat{F}=0,$ (41)
that is,
$\delta_{\tilde{\eta}}\tilde{F}+\iota_{b^{\sharp}}\tilde{F}=0,\qquad d\tilde{F}+b\wedge\tilde{F}-b\wedge F_{d}=0,\qquad\delta_{\tilde{\eta}}F_{d}-\iota_{b^{\sharp}}\tilde{F}=0,\qquad dF_{d}=0.$ (42)
Three-vector source-free formulation. For our next usage, and because it is more convenient, we obtain as well, the conventional three-vector formalism of Maxwell equations. To this end, we choose global inertial coordinates $(t,\mathbf{x})$ on $(M,\eta)$ and decompose
$\tilde{A}_{\mu}=(\tilde{\phi},\tilde{A}),\qquad A_{d,\mu}=(\phi_{d},A_{d}),\qquad b_{\mu}=(\partial_{t}\lambda,\nabla\lambda).$
(Spatial vectors are written without arrows.) Then, by definition of the generalized fields
$\boxed{\ \begin{aligned} \tilde{E}&:=-\partial_{t}\tilde{A}-\nabla\tilde{\phi}\;-\;(\partial_{t}\lambda)\,(\tilde{A}-A_{d})\;+\;(\tilde{\phi}-\phi_{d})\,\nabla\lambda,\\[2.0pt] \tilde{B}&:=\nabla\times\tilde{A}\;+\;(\nabla\lambda)\times(\tilde{A}-A_{d}),\end{aligned}}$ (43)
and, for the reference slot, $E_{d}:=-\partial_{t}A_{d}-\nabla\phi_{d}$, $B_{d}:=\nabla\times A_{d}$, the homogeneous generalized Maxwell equations (42) translate to
$\nabla\cdot\tilde{E}\;+\;(\nabla\lambda)\cdot(\tilde{E}-E_{d})$ $=0,$ (44)
$\partial_{t}\tilde{E}\;-\;\nabla\times\tilde{B}\;+\;(\partial_{t}\lambda)\,(\tilde{E}-E_{d})\;-\;(\nabla\lambda)\times(\tilde{B}-B_{d})$ $=0,$ (45)
$\partial_{t}\tilde{B}\;+\;\nabla\times\tilde{E}\;+\;(\partial_{t}\lambda)\,(\tilde{B}-B_{d})\;+\;(\nabla\lambda)\times(\tilde{E}-E_{d})$ $=0,$ (46)
$\nabla\cdot\tilde{B}\;+\;(\nabla\lambda)\cdot(\tilde{B}-B_{d})$ $=0,$ (47)
which reduce to the standard sourceless Maxwell equations when $\lambda$ is constant (so $b=d\lambda=0$) and $A_{d}$ is fixed (hence $E_{d}=B_{d}=0$).
Remarks.
  • (i)
    Gauge symmetry: for any scalar doublet $\hat{\chi}=(\chi,\chi_{d})$, the transformation $\hat{A}\mapsto\hat{A}+\hat{d}\,\hat{\chi}$ preserves $\hat{F}$ and $S_{\rm HGT}$, inducing the usual shift $\tilde{A}\mapsto\tilde{A}+d\chi$ in the upper slot.
  • (ii)
[…]